Find the measure of an angle if it's measure is 9° less than eight times the measure of its complement. A) 79 B) 159 C)11 D)21
Answer by rameens_99(10)

angle = x
90-x = complement
x=8(90-x)-9
x=720-8x-9
x=711-8x
9x=711
_______________
x= 79

So the correct answer would be A) 79.



You're Welcome (-: